Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120908

Observers : Iair Arcavi, Yen-Chen. 
Data reduction team : K. Maguire, M. Sullivan, A. Gal-Yam, O. Yaron

Clear skies (finally). Taking the opportunity to do SOFI followups.
Hip63862 - STD
SN2011ja - BLP Followup [SOFI: Bg]
Seeing ~1.2". Good signal. Set before we could do Rg and imaging, will try imaging tomorrow if conditions are good (there will be no time for Rg).
Hip79270 - STD
SNhunt133 - BLP Followup [SOFI: Bg]
Set before we could do Rg and imaging, will try imaging tomorrow if conditions are good (there will be no time for Rg).
SN2009ip - BLP Followup [SOFI: JHKs]
Seeing ~1.2". Can see the SN separated from the star next to it (unlike on Sep 6...).
LSQ12dwl - PESSTO Followup [SOFI: JHKs]
Seeing ~1.1". Can see the SN. 
BD174708003 - Photometric STD
=== STOPPED FOR TOO 2:53-3:48 UT ===
Hip106951 - STD
LSQ12dwl - PESSTO Followup [SOFI: Bg Rg]
Not sure if the red side has much signal. 
Hip12852 - STD 
SN2012ec - BLP Followup [SOFI: JHKs Bg Rg]
Seeing 1.2"-1.4" - Nice signal in Bg, faint but seen in Rg.
Hip023719 - STD
R71 - BLP Followup [SOFI: JHKs Bg Rg]
Comments